0 0
Add a comment Improve this question Transcribed image text
Answer #1

1) F = \DeltaP A =

= (1.013-0.960)x(1.2x1.3) = 0.11448 kN

2) bouyant force > WM + WL

=> \rhowater x(n x pi R2L)g > 320 xg + \rholog x(n x pi R2L)g

=> 16.579 n > 320

=> nmin = 19.3 = 20 logs

3) Q = Q(-12 to 0) + Q(s to L) +Q(0 to 5) = (3.2 x 2.108 x 12) kJ + (3.2 x 334) kJ + (3.2 x 4.186 x 5) kJ = 1216.7 kJ

4) When heat is transferred from one material to another it is called conduction. Metal is a good conductor of heat, while nonmetals like wood and plastic are poor conductors. The temperature of any metal object in any environment is about the same as the air that surrounds it. But in case of non metals there is a temperature difference.

When you touch a metal object that's surrounded by air cooler than your body, the object quickly conducts heat away from your fingertips, which makes them feel cold. The sensation travels to your brain, which perceives that the metal object is cold. But in case of plastic they don't adapt to the surroundings because of poor conductivity.

5) E= W = q x V = 5e x 20 V = 100 eV or 100x 1.6x10-19 J = 1.6x 10-17 J

6) F = (Kq1q3/r12 ) i + (Kq2q3/r22 )i      where i is a unit vector in positive X direction that is toeard -4 \muC

= (9x109 x 6 x10-6 x 3x10-6/0.052 ) -i + (9x109 x 4 x10-6 x 3x10-6/0.062 ) -i

= 94.8 N (-i)      or 94.8 N towards 6\muC

7) a)

The stream of water deviates a little towards the comb.

b) Water has no net electrical charge, but the negatively charged comb induces a complementary charge in the water. The electrons in the water can move somewhat. When the negatively charged comb comes near the water, it repels the electrons, so that the side of the water nearest the comb then has a positive charge. The attraction between this positive charge and the negatively charged comb results in a net force on the water, bending the stream.

8) v2 - u2 = 2aS => 0 - u2 = 2x-9.8 x 15

=> u = 17.146 m/s

therefore time to reach maximum height = (v-u) / a = (0-17.146)/-9.8 = 1.7496 s

a) so total time taken by ball to return = 2 x 1.7496 = 3.499 s

b) V before landing = velocity with which it was thrown = u = 17.146m/s